You need to raise the temperature of 55 g of water from 14.6 ◦C to 85◦C. How much heat is needed to accomplish this? The specific heat of water is 4180 J/kg · ◦ C. Answer in units of J
Answer #1

mass of water= 55g= 0.055kg

deltaT= 85-14.6=70.4 degrees

c=4180 J/kg · ◦ C

heat=Q= m c deltaT

Q= 0.055 * 4180 * (70.4)

Q=16184.96 j